OpenStudy (anonymous):

please help (: Find the x-, y-, and z-intercepts for the equation. Enter your answers separated by commas; example: -2, 3.5, 4 3x – 5y + 10z = 15

OpenStudy (anonymous):

Let's walk through this step by step. This is a 3-Dimensional graphical approach that will not, in the end, yield you answers that are integers, rather comparative variate's. Do you understand that means that instead of answers like -2, 3.5, and 4 - we will get answers like x = 2y + 1/z, y = 2/3x + z, etc. or something, though those are not even close to the solutions to this problem. Does this make sense so far? @Lovee_Stasiaa 

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):